# Hallucination to Truth: A Review of Fact-Checking and Factuality Evaluation in Large Language Models

### v1

@@ 4.2.2  Mitigation Strategies for LLM Hallucinations
Reinforcement learning has been used in the RAG process as well. For instance, Zhang et al. [51] investigated reinforcement learning to optimize retrieval processes; [34] propose an iterative retrieval and verification mechanism to refine accuracy further; Singhal et at. [53] introduce a RAG-based fact-checking system where the core pipeline retrieves top-3 documents with FAISS, extracts evidence, then uses the evidence in classification combining In-Context Learning (ICL) capabilities of multiple LLMs, achieving a 22% gain on Averitec dataset; Tran et al. [30] introduce Fact-Check-Then-RAG which improves LLM factual accuracy by evaluating each fact with SAFE [75], retrieving relevant data using ColBERT from MedRAG for failed checks, and re-generating responses via a RAG-enhanced prompt; Zhao et al. [14] introduced a retrieval-augmented and multi-agent fact-checking system which dynamically selects reasoning tools and external evidence to handle diverse multi-hop verification tasks.
@@
To enhance out-of-context (OOC) detection, Qi et al. [31] integrate Google’s Entity Detection API for visual grounding and performs external verification through LLMs to verify news captions against evidence retrieved from reverse image searches; [20] introduced SELF-CHECKER which verifies input by extracting simple claims, generating search queries on various external knowledge sources e.g., Bing Search API, retrieving evidence from knowledge sources e.g., Wikipedia, Reddit messages, and predicting each claim’s veracity based on its selected evidence sentences; Giarelis et al. [18] proposed a unified LLM-KG framework for fact-checking which retrieves relevant facts from Knowledge Graphs (KG) and injects them into the LLM prompt for reponse generation; Ma et al. [16] introduces Logical and Causal fact-checking method (LoCal), a LLM-driven multi-agent framework that breakdown complex claims, resolve them through specialized reasoning, and validate consistency using logical and counterfactual evaluators in an iterative process. Zhao et al. [49] introduced MEDICO where the system retrieves evidence from various sources search engine, knowledge base (KB), knowledge graph (KG) and user files, then re-ranks and fuses it by concatenation or Llama3-8B-based summarization; Li et al. [72] introduced FactAgent that uses LLM’s internal knowledge (i.e., Phrase, Language, Commonsense, Standing tools) and another that integrate external knowledge tools (i.e., URL and Search tools - SerpAPI); Ghosh et al. [17] introduced LLMQuery that evaluates LLMs’ logical consistency in fact-checking by retrieving subgraphs from KGs using BFS or ANN-based vector embedding methods, ensuring concise and relevant context is fed to the model; and finally, Tang et al. [10] introduce an efficient method specifically tailored for evaluating LLM-generated claims against grounding documents, forming a core component of comprehensive RAG evaluation.
@@
Adversarial Tuning. Adversarial training presents LLMs with specifically designed examples intended to uncover hallucinations or factual errors, thus training the models to recognize and handle these challenging inputs accurately and improve their robustness against generating misinformation. The study by Leippold et al. introduces CLIMINATOR, an acronym for CLImate Mediator for INformed Analysis and Transparent Objective Reasoning, an AI-based tool utilizing a Mediator and adversarial Advocate framework to automate climate claim verification by simulating structured debates, including climate denial perspectives, iteratively reconciling diverse viewpoints to consistently converge towards scientific consensus and thus improving accuracy and reliability [76].
@@
Automated Feedback Mechanisms and Self-Correction. Incorporating automated feedback loops and enabling LLMs to self-critique and correct their outputs are emerging as powerful strategies, exemplified by several key approaches: Peng et al. [15] proposed an automated feedback mechanism, LLM-AUGMENTER which substantially reduces ChatGPT’s hallucinations without sacrificing the fluency and informativeness of its responses by iteratively revising LLM prompts to improve model responses using feedback generated by utility functions, e.g., the factuality score of a LLM-generated response. Ma et al. [16] utilized two evaluating agents that iteratively reject or accept solutions and trigger new decomposition or reasoning rounds until consistency is achieved. Additionally, iterative refinement processes, such as those demonstrated by Xie et al. [34], continuously check and improve model outputs through multiple verification cycles; Tran et al. [30] introduced LEAF, a self-training loop that utilizes fact-check scores as automated feedback.
@@
Furthermore, the approach of uncertainty quantification proposes leveraging token-level uncertainty measures to detect potential hallucinations and trigger additional verification steps, thereby enhancing overall reliability and accuracy [77]. Ge et al. [43] also showed in their framework that LLMs leverage object-detector and VQA results to automatically mitigate hallucinated contents and fact-check proposed captions. In the work of Zhao et al. [49], it iteratively corrects only hallucinated parts in generated content using Chain-of-Thought (CoT) prompting, while enforcing minimal edits via Levenshtein-based preservation scoring.
@@
Hybrid Approaches and Multi-Agent Systems. Combining multiple strategies or employing multi-agent architectures, wherein different LLM agents handle specialized sub-tasks within the fact-checking process, has emerged as a growing trend, exemplified by hierarchical prompting and planning methods such as the work by Zhang et al. [21] systematically guide models through complex claim verification, and Zhao et al. [14] utilize LLMs for structured planning and reasoning tasks.
@@
Additionally, multi-agent systems have gained attention, as seen in studies like LoCal [16], employing multiple specialized LLM agents (decomposer + reasoner + two evaluators) to address logical and causal dimensions of fact-checking, and further explored by Kupershtein et al. [13] and Li et al. [72], both of which investigate the capabilities of agent-based frameworks specifically for fake news detection. Li et al. [72] also introduced FactAgent that behaves in an agentic manner, emulating human expert behavior utilizing several tools (Phrase, Language, Commonsense, Standing, URL, and Search tools) around a single LLM.
@@
Moreover, the integration of structured knowledge sources is further explored, advocating the combination of LLMs with knowledge graphs to leverage structured factual information [18]. Additionally, Hu et al. [12] designed a novel approach, ARG and its distilled version ARG-D, that complements small and large LMs by selectively acquiring insights from LLM-generated rationales for SLMs; this combination of LLM+SLM has shown superiority over existing SLM/LLM-only methods. By introducing a hybrid KG retrieval + LLM generation approach and supervised fine-tuning, Ghosh et al. [17] improved the logical consistency of LLMs on the complex fact-checking task. Similarly, Jing et al. [50] combined rubric-prompted LLM judges and an NLI cross-encoder (HHEM) in the same framework.
@@ 4.2.3  Recent Innovations for Reducing Hallucinations and Improving Factuality
Beyond the above core strategies, recent studies have further tried to address hallucination issues and ensure LLMs remain faithful to facts. These methods range from smarter prompting techniques and multi-step reasoning procedures to incorporating multiple modalities, to building self-checking mechanisms and uncertainty estimates into LLM responses. We highlight several promising directions below, along with their advances in architecture, evaluation, and practical deployment.
@@
Multimodal Fact-Checking. Misinformation is increasingly multimodal, combining text, images, and videos, thus addressing hallucinations and ensuring factuality in LLMs processing such data has become crucial. Efforts to integrate visual and textual evidence include the study by Cao et al. [36], which employs graph attention networks to consolidate multimodal knowledge for verifying claims, and Ge et al. [43], focusing specifically on accurate captioning of images utilizing visual fact-checking through object detection and VQA models. Furthermore, Sharma et al. [78] evaluate the visual grounding capabilities inherent in language models, while [31] introduced a multimodal LLM-SNIFFER which analyzes both the consistency of the image-text content and the claim-evidence relevance using InstructBLIP and GPT-4V. Practical applications are explored in [57], which investigates real-world deployment scenarios of multimodal LLMs, and finally, Khaliq et al. [22] specifically applies retrieval-augmented reasoning (ToRAG, CoRAG) to tackle multi-modal claims within political context by extracting both textual and image content, retrieving external information, and reasoning subsequent questions to be answered based on prior evidence and achieved a weighted F1-score of 0.85, surpassing a baseline reasoning method by 0.14 points.
Multilingual Fact-Checking. Misinformation transcends language barriers, necessitating fact-checking capabilities across multiple languages. LLMs offer significant potential in this area, but ensuring factual consistency across languages remains a considerable challenge. For instance, Quelle et al. [48] found that fact-checking accuracy varied across languages, with translated English prompts often achieving higher accuracy than original non-English ones in terms of multilingual fact-checking, despite claims involving non-English sources. Additionally, due to skewed training data and non-standardized fact-checks, LLMs perform better with English-translated prompts, revealing language bias in multilingual fact verification.
@@
In medical contexts, dedicated efforts include [28] introduced CliniFact and when evaluating LLMs against that, BioBERT achieved 80.2% accuracy, outperforming generative counterparts, such as Llama3-70B’s 53.6%, with statistical significance (p < 0.001), developing explainable reasoning systems [38], and detecting and correcting hallucinations by integrating multi-source evidence [49]. Similarly, Tans et al. [30] proposed LEAF, which tailors towards the medical domain as it uses the MedRAG corpus.
@@
Fact-checking is also investigated across various domains, including travel, climate, news information, and claim matching. For example, in the travel domain, Jing et al. [50] used four industry datasets containing chats, reviews, and property information for fact-checking. Political claim verification is explored through studies assessing LLM reliability [54] and multimodal retrieval-augmented reasoning systems [22]. In climate science, specialized LLM-based tools address the complexity of climate-related claims [47]. News claims and general misinformation are broadly examined through hierarchical prompting methods in [21], [12] investigated the potential of LLMs in fake news detection using Chinese dataset Weibo21 [81] and GossipCop [82], [55] focuses on news-article veracity: FA-KES Syrian War corpus [83] and EUvsDisinfo [84] pro-Kremlin corpus, and platforms enabling customized fact-checking system development [68]. [31] trained their SNIFFER model on news domain using the NewsCLIPpings [74] dataset. Additionally, claim-matching methods utilizing LLMs for fact-checking [80, 47, 41] and verification approaches for complex claims [37] further underscore the depth and breadth of ongoing domain-specific fact-checking research.
@@
Enhancing Explainability and Trust. Beyond mere accuracy, the ability of LLM-based fact-checking systems to provide explanations and foster trust is increasingly recognized as vital. Citations and provenance play a central role in this, with [56] highlighting the importance of referencing sources to build user confidence, while [39] emphasizes the need to trace the origin of generated content. In the framework proposed by Quelle et al. [48], agents explain their reasoning and cite the relevant sources. Additionally, Zhao et al. [14] enhanced fact-checking explainability by utilizing instruction-based LLMs and specialized agents to generate structured reasoning and justifications for sub-claims. Through CLIP-based similarity, ROUGE scores, response ratio analysis, and human evaluation, SNIFFER [31] demonstrates high accuracy and strong persuasive ability in explaining and detecting out-of-context misinformation, and transparency in medical contexts [38]. Credibility assessment frameworks offer structured approaches to evaluating trustworthiness [42]. Additionally, Ghosh et al. [17] investigate the logical soundness of model outputs, a key indicator of reliability. In another study, Leite et al. [55] simplify fact-checking by guiding LLMs to predict individual veracity signals, minimizing hallucinations, and enabling human reviewers to audit and control outputs, which enhances transparency. Giarelis et al. [18] stated that their LLM-KG framework improves transparency through factual context provided by Knowledge Graphs (KGs).
@@
Hierarchical Prompting and Multi-Step Reasoning. A key innovation involves prompting LLMs in a way that structures their reasoning process by decomposing complex fact-checking tasks into smaller, verifiable steps. The underlying idea is that human fact-checkers often break down a claim into sub-claims or evidence checks. LLMs can be prompted to emulate the process, reducing the risk of a single misstep leading to a hallucinated conclusion. For instance, Hierarchical Step-by-Step (HiSS) prompting directs the model to first separate a claim into several subclaims, then verify each subclaim one by one, before finalizing an overall verdict [21]. By forcing the model to focus on one piece of information at a time (in a chain-of-thought style), HiSS achieved superior fact verification performance and reduced hallucination on news datasets, even outperforming fully-supervised baselines. This demonstrates that prompting alone, if done cleverly, can induce the model to reason more carefully and factually reducing hallucination. In a study, Khaliq et al.[22] introduce Chain-of-RAG (sequential) and Tree-of-RAG (branch-and-eliminate hierarchy), which embody multi-step and hierarchical reasoning. Another example is the PACAR framework [14], which combines LLM-driven planning with customized action reasoning for claims. PACAR consists of multiple modules (a claim decomposer, a planner, an executor, and a verifier) that allow LLMs to plan a sequence of actions, such as performing a web search or a numerical calculation, and then verify the claim based on collected evidence. Using hierarchical prompting and a multi-step approach, which includes specialized skills like numerical reasoning and entity disambiguation, PACAR significantly outperformed baseline fact-checkers across three different domain datasets.

### v2

@@ 4.2.2  Mitigation Strategies for LLM Hallucinations
Reinforcement learning has also been used in the RAG process. For example, Zhang et al. [51] investigated reinforcement learning to optimize retrieval processes. [34] proposed an iterative retrieval and verification mechanism to refine accuracy further. Singhal et al. [53] introduced a RAG-based fact-checking system where the core pipeline retrieves top-3 documents with FAISS, extracts evidence, and then uses the evidence in classification, combining the In-Context Learning (ICL) capabilities of multiple LLMs, achieving a 22% gain on the Averitec dataset. Tran et al. [30] introduced Fact-Check-Then-RAG, which improves the factual accuracy of LLM by evaluating each fact with SAFE [74], retrieving relevant data using ColBERT from MedRAG for failed checks, and regenerating responses via a RAG-enhanced prompt. Zhao et al. [14] introduced a multi-agent and retrieval-augmented fact-checking system that dynamically selects reasoning tools and external evidence to handle diverse multi-hop verification tasks.
@@
To improve out-of-context (OOC) detection, Qi et al. [31] integrate Google’s Entity Detection API for visual grounding and perform external verification through LLM to verify news captions against evidence retrieved from reverse image searches. Li et al. [20] introduced SELF-CHECKER, which verifies input by extracting simple claims, generating search queries on various external knowledge sources, e.g., Bing Search API, retrieving evidence from knowledge sources, e.g., Wikipedia, Reddit messages, and predicting each claim’s veracity based on its selected evidence sentences. Giarelis et al. [18] proposed a unified LLM-KG framework for fact-checking, which retrieves relevant facts from Knowledge Graphs (KGs) and injects them into the LLM prompt for response generation. Ma et al. [16] introduced a Logical and Causal fact-checking method (LoCal), a LLM-driven multi-agent framework that breaks down complex claims, resolves them through specialized reasoning, and validates consistency using logical and counterfactual evaluators in an iterative process. Zhao et al. [49] introduced MEDICO, where the system retrieves evidence from various sources, search engines, knowledge bases (KB), KGs, and user files, then re-ranks and fuses it by concatenation or Llama3-8B-based summarization. Li et al. [71] introduced FactAgent that uses LLM’s internal knowledge (i.e., Phrase, Language, Commonsense, Standing tools) and another that integrates external knowledge tools (i.e., URL and Search tools - SerpAPI). Ghosh et al. [17] introduced LLMQuery that evaluates LLMs’ logical consistency in fact-checking by retrieving subgraphs from KGs using BFS or ANN-based vector embedding methods, ensuring concise and relevant context is fed to the model. Finally, Tang et al. [10] introduce an efficient method specifically tailored to evaluate claims generated by LLM against grounding documents, which forms a core component of a comprehensive evaluation of RAG.
@@
Adversarial Tuning. Adversarial training presents LLMs with specifically designed examples intended to uncover hallucinations or factual errors, thus training the models to recognize and handle these challenging inputs accurately and improve their robustness against generating misinformation. The study by Leippold et al. [75] introduced CLIMINATOR, an acronym for CLImate Mediator for INformed Analysis and Transparent Objective Reasoning, and an AI-based tool utilizing a Mediator and adversarial Advocate framework to automate climate claim verification by simulating structured debates, including climate denial perspectives, iteratively reconciling diverse viewpoints to converge towards scientific consensus consistently and thus improving accuracy and reliability.
@@
Automated Feedback Mechanisms and Self-Correction. Incorporating automated feedback loops and enabling LLMs to self-critique and correct their outputs are emerging as powerful strategies, exemplified by several key approaches. Peng et al. [15] proposed an automated feedback mechanism, LLM-AUGMENTER, which substantially reduces ChatGPT’s hallucinations without sacrificing the fluency and informativeness of its responses by iteratively revising LLM prompts to improve model responses using feedback generated by utility functions, e.g., the factuality score of a LLM-generated response. Ma et al. [16] utilized two evaluating agents that iteratively reject or accept solutions and trigger new decomposition or reasoning rounds until consistency is achieved. Additionally, iterative refinement processes, such as those demonstrated by Xie et al. [34], continuously check and improve model outputs through multiple verification cycles; Tran et al. [30] introduced LEAF, a self-training loop that utilizes fact-check scores as automated feedback.
@@
Furthermore, the uncertainty quantification approach proposes using token-level uncertainty measures to detect potential hallucinations and trigger additional verification steps, thereby enhancing overall reliability and accuracy [76]. Ge et al. [43] also showed in their framework that LLMs utilize object-detector and VQA results to mitigate hallucinated contents automatically and fact-check proposed captions. In the work of Zhao et al. [49], it iteratively corrects only hallucinated parts in generated content using Chain-of-Thought (CoT) prompting, while enforcing minimal edits via Levenshtein-based preservation scoring.
@@
Hybrid Approaches and Multi-Agent Systems. Combining multiple strategies or employing multi-agent architectures, wherein different LLM agents handle specialized sub-tasks within the fact-checking process, has emerged as a growing trend, exemplified by hierarchical prompting and planning methods such as the work by Zhang et al. [21] systematically guided models through complex claim verification, and Zhao et al. [14] utilized LLMs for structured planning and reasoning tasks.
@@
Additionally, multi-agent systems have gained attention, as seen in studies like LoCal [16], employing multiple specialized LLM agents (decomposer + reasoner + two evaluators) to address logical and causal dimensions of fact-checking. Further explored by Kupershtein et al. [13] and Li et al. [71], both of which investigate the capabilities of agent-based frameworks specifically for fake news detection. Li et al. [71] also introduced FactAgent that behaves in an agentic manner, emulating human expert behavior utilizing several tools (Phrase, Language, Commonsense, Standing, URL, and Search tools) around a single LLM.
@@
Moreover, the integration of structured knowledge sources is further explored, advocating the combination of LLMs with knowledge graphs to comprise structured factual information [18]. Additionally, Hu et al. [12] designed a novel approach, ARG and its distilled version ARG-D, that complements small and large LMMs by selectively acquiring insights from LLM-generated rationales for SLMs. This combination of LLM+SLM has shown superiority over existing SLM/LLM-only methods. By introducing a hybrid KG retrieval + LLM generation approach and supervised fine-tuning, Ghosh et al. [17] improved the logical consistency of LLMs on the complex fact-checking task. Similarly, Jing et al. [50] combined rubric-prompted LLM judges and an NLI cross-encoder (HHEM) in the same framework.
@@ 4.2.3  Recent Innovations for Reducing Hallucinations and Improving Factuality
Beyond the above core strategies, recent studies have further tried to address hallucination issues and ensure LLMs remain faithful to facts. These methods range from innovative prompting techniques and multi-step reasoning procedures to incorporating multiple modalities, as well as building self-checking mechanisms and uncertainty estimates into LLM responses. We highlight several promising directions below, along with their advances in architecture, evaluation, and practical deployment.
@@
Multimodal Fact-Checking. Misinformation is increasingly multimodal, combining text, images, and videos, thus addressing hallucinations and ensuring factuality in LLMs processing such data has become crucial. Efforts to integrate visual and textual evidence include the study by Cao et al. [36], which employed graph attention networks to consolidate multimodal knowledge for verifying claims, and Ge et al. [43], focusing specifically on accurate captioning of images utilizing visual fact-checking through object detection and VQA models. Then, to address multimodal fact-checking with explanation generation, Yao et al. [77] presented Mocheg, a large-scale benchmark dataset, and an initial demonstration of existing methods’ performance, which generated unsatisfactory results. Subsequently, Papadopoulos et al. [78] proposed the RED-DOT model, integrating a module called "relevant evidence detection" (RED) that can evaluate the relevancy of each piece of evidence and achieves up to 33.7% accuracy gain on the VERITE benchmark. Furthermore, Sharma et al. [79] evaluated the visual grounding capabilities inherent in language models, while [31] introduced a multimodal LLM-SNIFFER which analyzes both the consistency of the image-text content and the claim-evidence relevance using InstructBLIP and GPT-4V. Fake news detection is an emerging and vital domain addressed by Wang et al. [80], where they proposed a hybrid model called FND-LLM that combines SLMs and LLMs. This results in significant accuracy improvements of 0.7%, 6.8%, and 1.3% on the Weibo [90], Gossipcop [12], and Politifact datasets. Recently, Kakizaki et al. [81] introduced a system called Multimodal Automated Fact-checking via Textualization (MAFT), that textualizes content such as images, video, and audio for LLM analysis, which generates comprehensive and interpretable reports that explain the verification steps to combat misinformation. Practical applications are explored in [57], which investigates real-world deployment scenarios of multimodal LLMs, and finally, Khaliq et al. [22] specifically applied retrieval-augmented reasoning (ToRAG, CoRAG) to tackle multimodal claims within political context by extracting both textual and image content, retrieving external information, and reasoning subsequent questions to be answered based on prior evidence and achieved a weighted F1-score of 0.85, surpassing a baseline reasoning method by 0.14 points.
Multilingual Fact-Checking. Misinformation transcends language barriers, necessitating fact-checking capabilities across multiple languages. Zhang et al. [82] in their work conducted a systematic quantitative and qualitative analysis of the multilingual abilities of LLMs by introducing a novel prompt back-translation method. Their result reveals that LLMs excel in transferring learned knowledge across different languages, producing relatively consistent results in translation-equivariant tasks but struggle with translation-variant tasks, requiring careful human evaluation. Therefore, assessing the cross-lingual factual consistency of LLMs should be a primary concern in current research. For instance, Shafayat et al. [83] proposed a pipeline called Multi-FAct to evaluate the factuality of long-form multilingual LLM generations. This pipeline demonstrates the feasibility of adapting FActScore with non-English resources as a knowledge source, highlights their potential and limitations, and investigates cultural and geographic biases in LLMs. Quelle et al. [48] found that fact-checking accuracy varied across languages, with translated English prompts often achieving higher accuracy than original non-English ones in terms of multilingual fact-checking, despite claims involving non-English sources. Additionally, due to skewed training data and non-standardized fact-checks, LLMs perform better with English-translated prompts, revealing language bias in multilingual fact verification. We also found a unique and noteworthy work by Siino et al. [84] which introduced a comparative evaluation of state-of-the-art models, including LLMs for detecting fake news. Utilizing the multilingual Profiling Fake News Spreaders on Twitter (PFNSoT) dataset, which includes English and Spanish tweets, they showed that large, pre-trained Transformer models such as RoBERTa, DistilBERT, BERT, XLNet, ELECTRA, and Longformer are not necessarily the optimal solution for the classification task; instead, a convolutional neural network (CNN) is enough to outperform those. Their claim was validated by their proposed CNN model, demonstrating a binary accuracy of 71.5% for the English dataset and 81.5% for the Spanish dataset, where the highest performing transformer models, RoBERTa, only achieved an accuracy of 69.5% in English and ELECTRA achieved 76% in Spanish. Notably, in recent years, Shcharbakova et al. [85] demonstrated a comprehensive benchmark of language models for multilingual fact verification. Their experiments showed that the smaller encoder-based language model XLM-R significantly outperformed a much larger decoder-only LLM, achieving a macro-F1 score of 57.7%, which represents an improvement of approximately 15.8% over the previous state-of-the-art results. They suggested specialized SLMs can be more effective than general-purpose LLMs but also mentioned systematic difficulties in utilizing evidence effectively, alongside notable biases toward frequent categories within imbalanced data settings. In multilingual fact-checking, we found a distinctive and noteworthy work, performed by Jannah et al. [86], where they assessed the symptom detection capabilities of LLMs in social media posts across seven languages. Their experiment on zero-shot multi-label symptom classification includes large- and small-parameter models from three leading LLM providers: OpenAI, Google Gemini, and Mistral AI. They showed significant performance disparities, with LLMs performing better in European languages than in under-resourced Asian languages, and tending to over-predict specific respiratory illnesses like influenza.
@@
In medical contexts, dedicated efforts include [28], introduced CliniFact and when evaluating LLMs against that, BioBERT achieved 80.2% accuracy, outperforming generative counterparts, such as Llama3-70B’s 53.6%, with statistical significance (p < 0.001), developing explainable reasoning systems [38], and detecting and correcting hallucinations by integrating multi-source evidence [49]. Similarly, Tans et al. [30] proposed LEAF, which is tailored towards the medical domain as it uses the MedRAG corpus.
@@
Fact-checking is also investigated across various domains, including travel, climate, news information, and claim matching. For example, in the travel domain, Jing et al. [50] used four industry datasets containing chats, reviews, and property information for fact-checking. Political claim verification is explored through studies assessing LLMs’ reliability [54] and multimodal retrieval-augmented reasoning systems [22]. In climate science, specialized LLM-based tools address the complexity of climate-related claims [47]. News claims and general misinformation are broadly examined through hierarchical prompting methods in [21], [12]. Other studies investigated the potential of LLMs for fake news detection using datasets such as Weibo21 [91] and GossipCop [92]. Research on news-article veracity has focused on the FA-KES Syrian War corpus [93] and the EUvsDisinfo pro-Kremlin corpus [94], while some platforms enable the development of customized fact-checking systems [44]. [31] trained their SNIFFER model on news domain using the NewsCLIPpings [73] dataset. Additionally, claim-matching methods utilizing LLMs for fact-checking [41, 47, 88] and verification approaches for complex claims [37] further underscore the depth and breadth of ongoing domain-specific fact-checking research.
@@
Enhancing Explainability and Trust. Beyond mere accuracy, the ability of LLM-based fact-checking systems to provide explanations and foster trust is increasingly recognized as vital. Citations and provenance play a central role in this, with [56] highlighting the importance of referencing sources to build user confidence, while [39] emphasizes the need to trace the origin of generated content. In the framework proposed by Quelle et al. [48], agents explain their reasoning and cite the relevant sources. Zhao et al. [14] enhanced fact-checking explainability by utilizing instruction-based LLMs and specialized agents to generate structured reasoning and justifications for sub-claims. Through CLIP-based similarity, ROUGE scores, response ratio analysis, and human evaluation, SNIFFER [31] demonstrated high accuracy and strong persuasive ability in explaining and detecting out-of-context misinformation, and transparency in medical contexts [38]. Credibility assessment frameworks offer structured approaches to evaluate trustworthiness [42]. Additionally, Ghosh et al. [17] investigated the logical soundness of model outputs, a key indicator of reliability. In another study, Leite et al. [55] simplified fact-checking by guiding LLMs to predict individual veracity signals, minimizing hallucinations, and enabling human reviewers to audit and control outputs, which enhanced transparency. Giarelis et al. [18] stated that their LLM-KG framework improves transparency through factual context provided by KGs.
@@
Hierarchical Prompting and Multi-Step Reasoning. A key innovation involves prompting LLMs in a way that structures their reasoning process by decomposing complex fact-checking tasks into smaller, verifiable steps. The underlying idea is that human fact-checkers often break down a claim into sub-claims or evidence checks. LLMs can be prompted to emulate the process, reducing the risk of a single misstep leading to a hallucinated conclusion. For instance, HiSS prompting directs the model to first separate a claim into several subclaims, then verify each subclaim one by one, before finalizing an overall verdict [21]. By forcing the model to focus on one piece of information at a time (in a chain-of-thought style), HiSS achieved superior fact verification performance and reduced hallucination on news datasets, even outperforming fully-supervised baselines. Advance prompting like Hiss outperforms standard CoT Table 7. This demonstrates that well-designed prompting can guide the model to reason more carefully and factually reduce hallucination. In a study, Khaliq et al. [22] introduced Chain-of-RAG (sequential) and Tree-of-RAG (branch-and-eliminate hierarchy), which embodied multi-step and hierarchical reasoning. Another example is the PACAR framework [14], which combined LLM-driven planning with customized action reasoning for claims. PACAR consists of multiple modules (a claim decomposer, a planner, an executor, and a verifier) that allow LLMs to plan a sequence of actions, such as performing a web search or a numerical calculation, and then verify the claim based on collected evidence. Using hierarchical prompting and a multi-step approach, which includes specialized skills like numerical reasoning and entity disambiguation, PACAR significantly outperformed baseline fact-checkers across three different domain datasets.